# Benchmarking Web API Integration Code Generation

### v5

@@ V Conclusion and Future Work
Overall, our work underscores the need for integrating quality assurance techniques into AI-driven code generation workflows to ensure correctness and reliability. Our benchmark lays the foundation for future work to investigate the impact of different approaches—such as retrieval-augmented generation [49], fine-tuning, reasoning, feedback loops, and constrained decoding [50, 51]—on the correctness and specification-compliance of generated web API invocation code. While our evaluation pipeline is specialized on JavaScript and Axios, our dataset and methodology is generic and can be transferred to other languages and libraries.
Acknowledgments

### v6

@@ 8. Conclusion and Future Work
To address these challenges, we proposed the use of constrained decoding—a method that does not require costly retraining or additional context—for generating web API invocations. We developed a custom translation algorithm that derives constraints, represented as regular expressions, from OpenAPI specifications. This algorithm effectively manages the complexity and scale of real-world APIs, enabling the enforcement of compliance during code generation.
Our results demonstrate that constrained decoding, when equipped with appropriate constraints, can eliminate illegal and hallucinated endpoints and arguments by enforcing compliance with API specifications, substantially improving the accuracy and reliability of LLM-generated API code. Furthermore, our translation algorithm is implementation-agnostic and can be easily integrated as a plugin into other coding assistant systems to automatically enforce API specifications during development.
Overall, our work underscores the need for integrating quality assurance techniques into AI-driven code generation workflows to ensure correctness and reliability. While our focus has been on constrained decoding, our benchmark lays the foundation for future work to investigate the impact of further approaches—such as retrieval-augmented generation, fine-tuning, reasoning, or feedback loops—on the correctness and specification-compliance of generated web API invocation code. While our evaluation pipeline is specialized on JavaScript and Axios, our dataset and methodology is generic and can be transferred to other languages and libraries. Finally, we aim to extend constrained decoding to additional AI code generation tasks, broadening its impact and applicability.
Acknowledgements.